OpenStudy (anonymous):

Find the derivative of the function. f(x) = e^(5x + 7)

OpenStudy (anonymous):

I tried \[-e^{5x+7}\], but thats not it

OpenStudy (anonymous):

the derivative is f ' (x) = 5e^(5x+7)

OpenStudy (mimi_x3):

\[\large\frac{d}{dx} e^{ax+b} = e^{ax+b} *a\] \[\large\frac{d}{dx} e^{5x+7} = e^{5x+7} *5\]

OpenStudy (anonymous):

thanks guys